OpenStudy (tswiftcandle):

Write the equation of the line in Standard Form that passes through the points (3, 2) and (5, 8). Question 12 options: y = 7x + 3 -3x + y = 7 3x + y = -7 -3x + y = -7 @kalebj @linktra

OpenStudy (brainlesslonelygirl):

last option. do you want me to explain?

OpenStudy (mathmale):

First, a quiz for you! Write out the general form of the "Standard Equation of a straight line."

OpenStudy (tswiftcandle):

Yes please!! @brainless

OpenStudy (tswiftcandle):

@brainlesslonelygirl **

OpenStudy (brainlesslonelygirl):

well first you take out the gradient by the formula m=y2-y1/x2-x1 then the formula for a line passing a point is y-y1=m (x-x1) you know the gradient. take any of the two point and place them in the equation. and there you have your answer. try to attempt the question yourself now and see if you get it

OpenStudy (brainlesslonelygirl):

any of the two coordinates*